OpenStudy (anonymous):

There are 3 white marbles and 7 blue marbles in a bag. Jamie will randomly pick two marbles out of the bag without replacing the first marble chosen. What is the probability of Jamie's picking 2 white marbles? A) 1/15 B) 1/18 C) 1/20 D) 1/7

OpenStudy (anonymous):

A think

OpenStudy (anonymous):

THE PROBABILITY THAT HE WILL PICK THE FIRST WHITE MARBLE IS 3 /10 SINCE THE MARBLE IS NOT REPLACED THE TOTAL NUMBER OF MARBES NOW IS 9 AND REMAINING ARE 2 WHITE MARBLES PROBABILITY OF PICKING THE SECOND WHITE MARBLE = 2/9 so PROBABILITY OF PICKING TWO WHITE MARBLES = 3/10* 2/9 = 1/15

OpenStudy (anonymous):

A